# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ 5.1 Convergence in the range space
All the results presented above are concerned with the single step of (2) only. Now, we turn to the convergence analysis for the generated iterate sequence with the following assumption of existence of solutions.
Assumption 5
[Existence of solution to (1)] $\mathsf{zer}\mathcal{A}=\mathcal{A}^{-1}(0)\neq\varnothing$.
The following is an intermediate result from the proof of [22, Theorem 3.4], which is also crucial for our convergence analysis.
@@
Under Assumptions 1, 2 and 5, given $\mathcal{T}$ as (2), then
$\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A}).$
@@
(1) $x\in\mathsf{zer}\mathcal{A}\Longrightarrow 0\in\mathcal{A}x\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})x\Longrightarrow x\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x)\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. On the other hand, $x\in\mathsf{zer}\mathcal{A}\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Thus, we obtain $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})\subseteq\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$.
(2) Let $x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})x\Longrightarrow\mathcal{Q}x=\mathcal{Q}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})x=\mathcal{Q}\mathcal{T}x$. On the other hand, $\mathcal{T}=(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\Longrightarrow 0\in\mathcal{A}(\mathcal{T}x)+\mathcal{Q}\mathcal{T}x-\mathcal{Q}x$, $\forall x\in\mathsf{dom}\mathcal{T}$. Combining both above, we then have $x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x\in\mathcal{A}(\mathcal{T}x)\Longrightarrow\mathcal{T}x\subseteq\mathsf{zer}\mathcal{A}\Longrightarrow(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Since $x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)$, we have $x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Thus, $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\subseteq\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$.
Let us now discuss the convergence of (2). More concretely, we improve the weak convergence of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{k}\rightharpoonup\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$ established previously in [22, Theorem 3.4] by removing the finite dimensional restrict. For notational simplicity, we denote $x_{\mathsf{r}}^{k}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{k}$, $\forall k\in\mathbb{N}$.
@@
[Weak convergence in $\mathrm{ran}\mathcal{Q}$] Under Assumptions 1, 3 and 5, let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2). Then $x_{\mathsf{r}}^{k}\rightharpoonup\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$ for some $x^{\star}\in\mathsf{zer}\mathcal{A}$, as $k\rightarrow\infty$.
@@
Step-1: For every $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$, $\lim_{k\rightarrow\infty}$ $\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists.
Step-2: $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ has at least one weak sequential cluster point lying in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$;
@@
Step-1: First, we define $x_{\mathsf{r}}^{\star}\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. We also have $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ by Lemma 5.1, which implies that $\exists x^{\star}\in\mathsf{zer}\mathcal{A}$, such that $x_{\mathsf{r}}^{\star}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$. According to Fact 2, substituting $x=x_{\mathsf{r}}^{k}$ and $y=x_{\mathsf{r}}^{\star}$ into (7) yields
@@
Step-2: (16) implies the boundedness of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Indeed, from (16) and the closedness of $\mathsf{ran}\mathcal{Q}$, there exists $\alpha>0$, such that
@@
Thus, $\big{\|}x_{\mathsf{r}}^{k}\big{\|}\leq\big{\|}x_{\mathsf{r}}^{\star}\big{\|}+\frac{1}{\alpha}\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}:=\rho$, $\forall k\in\mathbb{N}$, and $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}\subseteq B(0,\rho)\cap\mathsf{ran}\mathcal{Q}$. $B(0,\rho)$ is weakly sequentially compact (by [4, Lemma 2.45]), and also is $B(0,\rho)\cap\mathsf{ran}\mathcal{Q}\subseteq B(0,\rho)$ (by [4, Lemma 1.34]). Then, $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ must possess at lease one weak sequential cluster point $v^{*}\in B(0,\rho)\cap\mathsf{ran}\mathcal{Q}$, i.e., there exists a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k\in\mathbb{N}}$ that weakly converges to $v^{*}$, as $k_{i}\rightarrow\infty$. We here need to show that $v^{*}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$, and more generally, every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ belongs to $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. To this end, summing up (16) from $k=0$ to $K-1$, and taking $K\rightarrow\infty$, we have
@@
This implies that $x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}=(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k})\rightarrow 0$, and also $(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k_{i}})\rightarrow 0$ as $k_{i}\rightarrow\infty$. Combining with $x_{\mathsf{r}}^{k_{i}}\rightharpoonup v^{*}$ and the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ (shown in Lemma 2.8), we obtain $(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(v^{*})=0\Longrightarrow v^{*}\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. Since $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ is an arbitrary weakly convergence subsequence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, we conclude that every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$, and furthermore, in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ as well, again by Lemma 5.1.
Step-3: We need to show that $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ cannot have two distinct weak sequential cluster point in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Indeed, let $v_{1}^{*},v_{2}^{*}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ be two cluster points of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Since $\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists as proved in Step-1, set $l_{1}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{1}^{*}\|_{\mathcal{Q}}$, and $l_{2}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{2}^{*}\|_{\mathcal{Q}}$. Take a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ weakly converging to $v_{1}^{*}$, as $k_{i}\rightarrow\infty$. From the identity:
@@
then, we deduce that $l_{1}^{2}-l_{2}^{2}=-\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Similarly, take a subsequence $\{x^{k_{j}}\}_{k_{j}\in\mathbb{N}}$ weakly converging to $v_{2}^{*}$, as $k_{j}\rightarrow\infty$, which yields that $l_{1}^{2}-l_{2}^{2}=\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Consequently, $\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}=0\Longrightarrow v_{1}^{*}=v_{2}^{*}$, since $v_{1}^{*},v_{2}^{*}\in\mathsf{ran}\mathcal{Q}$. This shows the uniqueness of the weak sequential cluster point.
Finally, combining the above 3 step and [4, Lemma 2.46], we conclude the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, and denote the weak limit as $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$.

### v4

@@ 5.1 Convergence in the range space
Now, we turn to the convergence analysis of (2). The following is an intermediate result from the proof of [33, Theorem 3.4], which is also crucial for our convergence analysis.
@@
Under Assumptions 1 and 2, given $\mathcal{T}$ as (2), then
$\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A}).$
@@
(1) $x\in\mathrm{zer}\mathcal{A}\Longrightarrow 0\in\mathcal{A}x\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})x\Longrightarrow x\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x=(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x)=(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x)\Longrightarrow\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x\in\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})$. On the other hand, $x\in\mathrm{zer}\mathcal{A}\Longrightarrow\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$. Thus, we obtain $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})\subseteq\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})$.
(2) Let $x\in\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x=(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})x\Longrightarrow\mathcal{Q}x=\mathcal{Q}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})x=\mathcal{Q}\mathcal{T}x$. On the other hand, $\mathcal{T}=(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\Longrightarrow 0\in\mathcal{A}(\mathcal{T}x)+\mathcal{Q}\mathcal{T}x-\mathcal{Q}x$, $\forall x\in\mathrm{dom}\mathcal{T}$. Combining both above, we then have $x\in\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x\in\mathcal{A}(\mathcal{T}x)\Longrightarrow\mathcal{T}x\subseteq\mathrm{zer}\mathcal{A}\Longrightarrow(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x)\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$. Since $x=(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x)$, we have $x\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$. Thus, $\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})\subseteq\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$.
Let us now discuss the convergence of (2). Theorem 5.3 establishes the weak convergence of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x^{k}\rightharpoonup\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x^{\star}$ without finite dimensional restriction used previously in [33, Theorem 3.4]. For notational simplicity, we denote $x_{\mathsf{r}}^{k}:=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x^{k}$, $\forall k\in\mathbb{N}$.
@@
[Weak convergence in $\mathrm{ran}\mathcal{Q}$] Under Assumptions 1 and 3, let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2). Then $x_{\mathsf{r}}^{k}\rightharpoonup\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x^{\star}$ for some $x^{\star}\in\mathrm{zer}\mathcal{A}$, as $k\rightarrow\infty$.
@@
Step-1: For every $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$, $\lim_{k\rightarrow\infty}$ $\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists.
Step-2: $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ has at least one weak sequential cluster point lying in $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$;
@@
Step-1: First, we define $x_{\mathsf{r}}^{\star}\in\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})$. We also have $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$ by Lemma 5.1, which implies that $\exists x^{\star}\in\mathrm{zer}\mathcal{A}$, such that $x_{\mathsf{r}}^{\star}=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x^{\star}$. According to Fact 2, substituting $x=x_{\mathsf{r}}^{k}$ and $y=x_{\mathsf{r}}^{\star}$ into (7) yields
@@
Step-2: (16) implies the boundedness of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Indeed, from (16) and the closedness of $\mathrm{ran}\mathcal{Q}$, there exists $\alpha>0$, such that
@@
Thus, $\big{\|}x_{\mathsf{r}}^{k}\big{\|}\leq\big{\|}x_{\mathsf{r}}^{\star}\big{\|}+\frac{1}{\alpha}\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}:=\rho$, $\forall k\in\mathbb{N}$, and $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}\subseteq B(0,\rho)\cap\mathrm{ran}\mathcal{Q}$. $B(0,\rho)$ is weakly sequentially compact (by [5, Lemma 2.45]), and also is $B(0,\rho)\cap\mathrm{ran}\mathcal{Q}\subseteq B(0,\rho)$ (by [5, Lemma 1.34]). Then, $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ must possess at least one weak sequential cluster point $v^{*}\in B(0,\rho)\cap\mathrm{ran}\mathcal{Q}$, i.e., there exists a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k\in\mathbb{N}}$ that weakly converges to $v^{*}$, as $k_{i}\rightarrow\infty$. We here need to show that $v^{*}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$, and more generally, every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ belongs to $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$. To this end, summing up (16) from $k=0$ to $K-1$, and taking $K\rightarrow\infty$, we have
@@
This implies that $x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}=(\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k})\rightarrow 0$, and also $(\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k_{i}})\rightarrow 0$ as $k_{i}\rightarrow\infty$. Combining with $x_{\mathsf{r}}^{k_{i}}\rightharpoonup v^{*}\in\mathrm{ran}\mathcal{Q}$ and the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}$ (shown in Lemma 2.9), we obtain $(\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}})(v^{*})=(\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(v^{*})=0\Longrightarrow v^{*}\in\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})$. Since $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ is an arbitrary weakly convergent subsequence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, we conclude that every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})$, and furthermore, in $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$ as well, again by Lemma 5.1.
Step-3: We need to show that $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ cannot have two distinct weak sequential cluster point in $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$. Indeed, let $v_{1}^{*},v_{2}^{*}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$ be two cluster points of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Since $\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists as proved in Step-1, set $l_{1}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{1}^{*}\|_{\mathcal{Q}}$, and $l_{2}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{2}^{*}\|_{\mathcal{Q}}$. Take a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ weakly converging to $v_{1}^{*}$, as $k_{i}\rightarrow\infty$. From the identity:
@@
then, we deduce that $l_{1}^{2}-l_{2}^{2}=-\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Similarly, take a subsequence $\{x^{k_{j}}\}_{k_{j}\in\mathbb{N}}$ weakly converging to $v_{2}^{*}$, as $k_{j}\rightarrow\infty$, which yields that $l_{1}^{2}-l_{2}^{2}=\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Consequently, $\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}=0\Longrightarrow v_{1}^{*}=v_{2}^{*}$, since $v_{1}^{*},v_{2}^{*}\in\mathrm{ran}\mathcal{Q}$. This shows the uniqueness of the weak sequential cluster point.
Finally, combining the above 3 step and [5, Lemma 2.46], we conclude the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, and denote the weak limit as $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A})$.